OpenStudy (anonymous):

What is the solution set for |–2x + 8| = 12 x = –2 and x = 10 x = –2 and x = –10 x = 2 and x = 10 x = 2 and x = –10

OpenStudy (anonymous):

i think its C

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

umm it's definitely a) because 4+8 is 12 and -20+8 is -12 but since it's absolute value it's just 12

OpenStudy (anonymous):

yeah its a. i had to look at it again.

OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (1, –3) and has a slope of 2? 2x – y = 5 2x – y = –5 2x + y = 5 2x + y = –5